OpenStudy (gayaznpanda94):

prove that x/e^x + x/2 = x(e^x+1)/2(e^x-1)

OpenStudy (gayaznpanda94):

* x/e^x +1 +X/2

OpenStudy (gayaznpanda94):

x/(e^x+1) +x/2

OpenStudy (phi):

Is the problem \[\frac{x}{e^x+1}+\frac{x}{2}\]?

OpenStudy (gayaznpanda94):

yes

OpenStudy (phi):

you can put them over a common denominator and add \[\frac{2x}{2(e^x+1)}+\frac{x(e^x+1)}{2(e^x+1)} \]

OpenStudy (phi):

If you do, you will see that you get \[\frac{x}{2}\frac{e^x+3}{e^x+1}\]

OpenStudy (gayaznpanda94):

i figured it out thank you tho

OpenStudy (phi):

How? because the answer does not match your question x(e^x+1)/2(e^x-1)

OpenStudy (gayaznpanda94):

it should be x(e^x-1) i accidentally typed it wrong

OpenStudy (phi):

OK, that makes sense. Just wondering.